Using the classical method, solve

(D? + 2D y(t) = (D+1) f(t)

if the initial conditions are y(0) = 2 , Y(0) = 1 , and if the input if f(t)=u(t)

Do not use convolution.
Answer #1

* Given that differential Equation (D?+20) y(t) = (0+1) FC) f (t) = UCt) 410)=2, 4(0)=1 i we have to find complete solution 1given another initial conditim Ý (O) = 1 2Cqe2t to It) = at to Ý (0) = 2 C eo 1 = 2 = 1/2 by putting value of Ca in Eqn (1) G